# web_appendix

### v4

@@ 2.3 Building-block identification result
If Assumption 1 holds, then for all $(d_{1},d_{2})\in\mathcal{D}_{1}\times\mathcal{D}_{2}$ such that $d_{1}\neq d_{2}$ and $P(S|D_{1}=d_{1})<1$,
@@
Note that in a canonical DID design where $\mathcal{D}_{1}=0$ and $\mathcal{D}_{2}\in\{0,1\}$, Lemma 1 only applies to $(d_{1},d_{2})=(0,1)$, $\text{TE}(0,1|0,1)$ reduces to the ATT, and the estimand reduces to the canonical DID estimand comparing the outcome evolutions of treated and untreated units. Note also that in a design where all units are untreated at period one, $d_{1}=0$, and
@@
an effect closely related to the $ATT(d|d)$ effect in Callaway et al. (2021). Thus, the effects we consider are extensions of those effects to applications with a treatment continuous at all periods.
Lemma 1 shows that under Assumption 1,
$(d_{1},d_{2})\mapsto\text{TE}(d_{1},d_{2}|d_{1},d_{2})$
is identified. Of course, one may be interested in other parameters, like
$(d,d^{\prime})\mapsto\text{TE}(d,d^{\prime}):=E\left(\frac{Y_{2}(d)-Y_{2}(d^{\prime})}{d-d^{\prime}}\right),$
a function which, unlike $(d_{1},d_{2})\mapsto\text{TE}(d_{1},d_{2}|d_{1},d_{2}),$ applies to the entire population rather than to specific subpopulation that depend on $(d_{1},d_{2})$. Alternatively, one could also be interested in the average marginal effect
$E\left(Y^{\prime}_{2}(D_{2})\right).$
What is the appeal of $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ with respect to those other parameters? Conditional on $D_{1}=d_{1},D_{2}=d_{2}$, $Y_{2}(d_{2})$ is observed, so estimating $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ only requires estimating $Y_{2}(d_{1})$, switchers’ counterfactual outcomes if their treatment had not changed. By definition, $Y_{1}(d_{1})$ is observed. If the data contains a third period 00 and the treatment of some units does not change from period 00 to $1$, then $Y_{0}(d_{1})$ is also observed for some switchers and stayers. Then, as explained in further details in Section 6.2, one can placebo-test Assumption 1, by comparing the outcome evolutions of switchers and stayers from period 00 to $1$. This shows that $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ is identified under a placebo-testable parallel-trends assumption. On the other hand, estimating $\text{TE}(d,d^{\prime})$ requires estimating, for most units, two unobserved counterfactual outcomes. This cannot be achieved under a placebo-testable assumption as we only observe one potential outcome at each date. When $D_{1}=0$, Callaway et al. (2021) propose a “strong parallel-trends” assumption under which the dose-response function, a parameter closely related to $\text{TE}(0,d^{\prime})$ are identified, but their “strong parallel-trends” assumption is not placebo testable. Similarly, estimating $E\left(\frac{Y_{2}(D_{2})-Y_{2}(d^{\prime})}{D_{2}-d^{\prime}}\right)$ requires estimating $Y_{2}(d^{\prime})$, which cannot be achieved under a placebo-testable assumption because $Y_{1}(d^{\prime})$ is not observed for all units. As $Y^{\prime}_{2}(D_{2})=\lim_{d^{\prime}\rightarrow D_{2}}\frac{Y_{2}(D_{2})-Y_{2}(d^{\prime})}{D_{2}-d^{\prime}}$, the same issue applies to $E\left(Y^{\prime}_{2}(D_{2})\right)$.
@@
Variability in $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ across values of $(d_{1},d_{2})$ conflates a dose-response relationship that may be of economic interest, and a selection bias due to the fact that units with different period one and two treatments may have heterogeneous treatment effects (Callaway et al., 2021). Moreover, Lemma 1 shows that estimating $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ requires estimating the values of two conditional expectations with respect to continuous variables, at points $D_{1}=d_{1},D_{2}=d_{2}$ and $D_{1}=d_{1}$. Unless one is willing to make parametric functional-form assumptions, the resulting estimator will converge at a slower rate than the standard $\sqrt{n}-$ parametric rate. For these two reasons, in this paper we focus on averages of the slopes $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$, that can be estimated non-parametrically at the standard $\sqrt{n}-$ parametric rate, and we do not focus on the function $(d_{1},d_{2})\mapsto\text{TE}(d_{1},d_{2}|d_{1},d_{2}).$
@@
Finally, our DID estimands compare switchers and stayers with the same period-one treatment. Instead, one could propose estimands comparing switchers and stayers, without conditioning on their period-one treatment. To recover the counterfactual outcome trend of a switcher going from two to five units of treatment, one could use a stayer with treatment equal to three at both dates. On top of Assumption 1, such estimands rest on two supplementary conditions:
@@
    For all $(d,d^{\prime})\in\mathcal{D}_{1}^{2}$, $E(\Delta Y(d))=E(\Delta Y(d^{\prime}))$.
(i) requires that all units experience the same evolution of their potential outcome with treatment $d$, while Assumption 1 only imposes that requirement for units with the same baseline treatment. Assumption 1 may be more plausible: units with the same period-one treatment may be more similar and more likely to be on parallel trends than units with different period-one treatments. (ii) requires that the trend affecting all potential outcomes be the same: to rationalize a DID estimand comparing a switcher going from two to five units of treatment to a stayer with treatment equal to three, $E(\Delta Y(2))$ and $E(\Delta Y(3))$ should be equal. Rearranging, (ii) is equivalent to
$E(Y_{2}(d)-Y_{2}(d^{\prime}))=E(Y_{1}(d)-Y_{1}(d^{\prime})):$ (1)
the treatment effect should be constant over time, a strong restriction on treatment effect heterogeneity. Assumption 1, on the other hand, does not impose any restriction on treatment effect heterogeneity, as it only restricts one potential outcome per unit.
@@ 3.1 Target parameter
Accordingly, under convexity one can use the AS to obtain lower bounds of the effect of changing the treatment from $D_{2}$ to larger values than $D_{2}$. For instance, in Fajgelbaum et al. (2020), one can use this strategy to derive a lower bound of the effect of increasing tariffs’ to even higher levels than those decided by the Trump administration. Under convexity, one can also use the AS to derive an upper bound of the effect of changing the treatment from $D_{1}$ to a lower value than $D_{1}$. And under concavity, one can derive an upper (resp. lower) bound of the effect of changing the treatment from $D_{2}$ (resp. $D_{1}$) to a larger (resp. lower) value. [^4] See D’Haultfœuille et al. (2023) for bounds of the same kind obtained under concavity or convexity. Importantly, the AS is identified even if those linearity or convexity/concavity conditions fail. But those non-placebo testable conditions are necessary to use the AS to identify or bound the effects of alternative policies.
@@ 3.2 Identification
To identify the AS, we use a DID estimand comparing switchers and stayers with the same period-one treatment. This requires that there be no value of the period-one treatment $D_{1}$ such that only switchers have that value, as stated formally below.
@@
(Support condition for AS identification) $P(S=1)>0$, $P(S=1|D_{1})<1$.
Assumption 4 implies that $P(S=0)>0$, meaning that there are stayers whose treatment does not change. While we assume that $D_{1}$ and $D_{2}$ are continuous, we also assume that the treatment is persistent, and thus $\Delta D$ has a mixed distribution with a mass point at zero.
To identify the AS, we also start by assuming that there are no quasi-stayers: the treatment of all switchers changes by at last $c$ from period one to two, for some strictly positive $c$.
@@
(No quasi-stayers) $\exists c>0$: $P(|\Delta D|>c|S=1)=1$.
We relax Assumption 5 just below.
@@
If Assumptions 1-5 hold,
$\delta_{1}$ $=E\left(\frac{\Delta Y-E(\Delta Y|D_{1},S=0)}{\Delta D}\middle|S=1\right).$
If there are quasi-stayers, the AS is still identified. For any $\eta>0$, let $S_{\eta}=1\{|\Delta D|>\eta\}$ be an indicator for switchers whose treatment changes by at least $\eta$ from period one to two.
@@
$\Delta Y-E(\Delta Y|D_{1},S=0)$
$=$ $Y_{2}(D_{2})-Y_{2}(D_{1})+\Delta Y(D_{1})-E(\Delta Y(D_{1})|D_{1},S=0)$
@@
Therefore, we need to trim quasi-stayers from the estimand in Theorem 1, and let the trimming go to 0, as in Graham and Powell (2012) who consider a related estimand with some quasi-stayers. Accordingly, with quasi-stayers the AS is irregularly identified by a limiting estimand.
@@ 3.3 Estimation and inference
With no quasi-stayers, $E\left((\Delta Y-E(\Delta Y|D_{1},S=0))/\Delta D\middle|S=1\right)$ can be estimated in three steps. First, one estimates $E(\Delta Y|D_{1},S=0)$ using a non-parametric regression of $\Delta Y_{i}$ on $D_{i,1}$ among stayers. Second, for each switcher, one computes $\widehat{E}(\Delta Y|D_{1}=D_{i,1},S=0)$, its predicted outcome evolution given its baseline treatment, according to the non-parametric regression estimated among stayers. Third, one lets
$\widehat{\delta}_{1}:=$ $\frac{1}{n_{s}}\sum_{i:S_{i}=1}\frac{\Delta Y_{i}-\widehat{E}(\Delta Y|D_{1}=D_{i,1},S=0)}{\Delta D_{i}},$
where $n_{s}=\#\{i:S_{i}=1\}$.
To estimate $E(\Delta Y|D_{1},S=0)$, we consider a series estimator based on polynomials in $D_{1}$, $(p_{k,K_{n}}(D_{1}))_{1\leq k\leq K_{n}}$. We make the following technical assumption.
@@
(Conditions for asymptotic normality of AS estimator)
  1. 1.
    $D_{1}$ is continuously distributed on a compact interval $I$, with $\inf_{d\in I}f_{D_{1}}(d)>0$.
  2. 2.
    $E[\Delta Y^{2}]<\infty$ and $d\mapsto E[\Delta Y^{2}|D_{1}=d]$ is bounded on $I$.
  3. 3.
    $P(S=1)>0$ and $\sup_{d\in I}P(S=1|D_{1}=d)<1$.
  4. 4.
    The functions $d\mapsto E[(1-S)\Delta Y|D_{1}=d]$, $d\mapsto E[S|D_{1}=d]$ and $d\mapsto E[S/\Delta D|D_{1}=d]$ are four times continuously differentiable.
  5. 5.
    The polynomials $d\mapsto p_{k,K_{n}}(d)$, $1\leq k\leq K_{n}$, are orthonormal on $I$ and $K_{n}^{12}/n\to+\infty$, $K_{n}^{7}/n\to 0$.
Point 3 is a slight reinforcement of Assumption 4. In Point 5, $K_{n}^{12}/n\to\infty$ requires that $K_{n}$, the order of the polynomial in $D_{1}$ we use to approximate $E(\Delta Y|D_{1},S=0)$, goes to +$\infty$+ ∞ when the sample size grows, thus ensuring that the bias of our series estimator of $E(\Delta Y|D_{1},S=0)$ tends to zero. $K_{n}^{7}/n\to 0$ ensures that $K_{n}$ does not go to infinity too fast, thus preventing overfitting.
@@
If Assumptions 1-3 and 5-6 hold,
$\sqrt{n}\left(\widehat{\delta}_{1}-\delta_{1}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{1})),$
where
$\psi_{1}$ $:=\frac{1}{E(S)}\left\{\left(\frac{S}{\Delta D}-E\left(\left.\frac{S}{\Delta D}\right|D_{1}\right)\frac{(1-S)}{E[1-S|D_{1}]}\right)[\Delta Y-E(\Delta Y|D_{1},S=0)]-\delta_{1}S\right\}.$
Theorem 3 shows that without quasi-stayers, the AS can be estimated at the $\sqrt{n}-$rate, and gives an expression of its estimator’s asymptotic variance. With quasi-stayers, we conjecture that the AS cannot be estimated at the $\sqrt{n}-$rate. This conjecture is based on a result from Graham and Powell (2012). Though their result applies to a broader class of estimands, it implies in particular that with quasi-stayers,
$\lim_{\eta\downarrow 0}E\left(\frac{\Delta Y-E(\Delta Y|S=0)}{\Delta D}\middle|S_{\eta}=1\right)$
cannot be estimated at a faster rate than $n^{1/3}$. The estimand in the previous display is closely related to our estimand
$\lim_{\eta\downarrow 0}E\left(\frac{\Delta Y-E(\Delta Y|D_{1},S=0)}{\Delta D}\middle|S_{\eta}=1\right)$
in Theorem 2, and is equal to it if $E(\Delta Y|D_{1},S=0)=E(\Delta Y|S=0)$. Then, even though the assumptions in Graham and Powell (2012) differ from ours, it seems reasonable to assume that their general conclusion still applies to our set-up: here as well, owing to $\delta_{1}$’s irregular identification, this parameter can probably not be estimated at the parametric $\sqrt{n}-$rate with quasi-stayers. This is one of the reasons that lead us to consider, in the next section, another target parameter that can be estimated at the parametric $\sqrt{n}-$rate with quasi-stayers.

### v5

@@ 2.3 Building-block identification result
If Assumption 1 holds, then $\forall$ $(d_{1},d_{2})\in\mathcal{D}_{1}\times\mathcal{D}_{2}$ such that $d_{1}\neq d_{2}$ and $P(S|D_{1}=d_{1})<1$,
@@
In a canonical DID design where $\mathcal{D}_{1}=\{0\}$ and $\mathcal{D}_{2}\in\{0,1\}$, Lemma 1 only applies to $(d_{1},d_{2})=(0,1)$, $\text{TE}(0,1|0,1)$ reduces to the average treatment on the treated (ATT), and the estimand reduces to the canonical DID estimand comparing the outcome evolutions of treated and untreated units. Also, if all units are untreated at period one ($\mathcal{D}_{1}=\{0\}$), we have
@@
an effect closely related to the $ATT(d|d)$ effect in Callaway et al. (2021). Thus, the effects we consider are extensions of those effects to applications with a treatment continuous in all periods.
@@
We do not propose estimators of $(d_{1},d_{2})\mapsto\text{TE}(d_{1},d_{2}|d_{1},d_{2})$, for two reasons. First, variability in $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ across values of $(d_{1},d_{2})$ conflates a dose-response relationship that may be of interest, and a selection bias, typically not of interest, due to the fact that units with different period one and two treatments may have different treatment effects (Callaway et al., 2021). Moreover, Lemma 1 shows that estimating $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ requires estimating two bivariate nonparametric regressions. Unless one is willing to make parametric functional-form assumptions, the resulting estimator converges at most at the $n^{1/3}$ rate (Stone, 1982), and may therefore be very imprecise, given that DID applications often have a relatively small number of units (e.g. the 50 US states in our empirical application). For these two reasons, in the next sections our target parameters will be two averages of the slopes $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$, which can be estimated nonparametrically at the $\sqrt{n}-$ parametric rate. In a recent paper, posterior to ours, Haddad et al. (2024) consider a parallel-trends assumption similar to Assumption 1, and propose an estimator of a functional parameter similar to $(d_{1},d_{2})\mapsto\text{TE}(d_{1},d_{2}|d_{1},d_{2})$.
Instead of averages of the slopes $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$, one may be interested in other parameters, like the average dose-slope function $(d,d^{\prime})\mapsto\text{TE}(d,d^{\prime}):=E\left((Y_{2}(d)-Y_{2}(d^{\prime}))/(d-d^{\prime})\right),$ or the average marginal effect $E\left(Y^{\prime}_{2}(D_{2})\right)$. However, as discussed in Section LABEL:appendixsec:alternatives of the web appendix, while averages of $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ can be identified under an assumption whose plausibility can be assessed via a pre-trends test, $\text{TE}(d,d^{\prime})$ and $E\left(Y^{\prime}_{2}(D_{2})\right)$ cannot be identified under assumptions whose plausibility can be assessed via a pre-trends test.
2.4 Alternative parallel-trends assumption
@@
The DID estimands in Lemma 1 compare switchers and stayers with the same period-one treatment. Instead, one could propose estimands comparing switchers and stayers, without conditioning on their period-one treatment. To recover the counterfactual outcome trend of a switcher going from two to five units of treatment, one could use a stayer with treatment equal to three at both dates. On top of Assumption 1, such estimands rest on two supplementary conditions:
@@
    $\forall$ $(d,d^{\prime})\in\mathcal{D}_{1}^{2}$, $E(\Delta Y(d))=E(\Delta Y(d^{\prime}))$.
(i) requires that all units experience the same evolution of their potential outcome with treatment $d$, while Assumption 1 only imposes that requirement for units with the same baseline treatment. Assumption 1 may be more plausible: units with the same period-one treatment may be more similar and more likely to be on parallel trends than units with different period-one treatments. (ii) requires that the trend affecting all potential outcomes be the same: to rationalize a DID estimand comparing a switcher going from two to five units of treatment to a stayer with treatment equal to three, $E(\Delta Y(2))$ and $E(\Delta Y(3))$ should be equal. Rearranging, (ii) is equivalent to $E(Y_{2}(d)-Y_{2}(d^{\prime}))=E(Y_{1}(d)-Y_{1}(d^{\prime}))$, so the treatment effect should be constant over time, a strong restriction on treatment effect heterogeneity. In contrast, our Assumption 1 does not restrict treatment effect heterogeneity.
@@ 3.1 Target parameter
Accordingly, under convexity one can use the AS to obtain lower bounds of the effect of changing the treatment from $D_{2}$ to larger values than $D_{2}$. For instance, in Fajgelbaum et al. (2020), one can use this strategy to derive a lower bound of the effect of increasing tariffs’ to even higher levels than those imposed in 2018 by the Trump administration. Under convexity, one can also use the AS to derive an upper bound of the effect of changing the treatment from $D_{1}$ to a lower value. And under concavity, one can derive an upper (resp. lower) bound of the effect of changing the treatment from $D_{2}$ (resp. $D_{1}$) to a larger (resp. lower) value. [^3] See D’Haultfœuille et al. (2023) for bounds of the same kind obtained under concavity or convexity. The AS is identified even if those linearity or convexity/concavity conditions fail, but those conditions are necessary to use the AS to identify or bound the effects of alternative policies.
@@ 3.2 Identification
The AS is identified under the two following assumptions on the design.
@@
(Overlap condition for AS identification) $P(S=1)>0$ and $P(S=0|D_{1})>\zeta$ for some $\zeta>0$.
@@
(No quasi-stayers) $P(|\Delta D|>\kappa|S=1)=1$ for some $\kappa>0$.
The estimand identifying the AS compares the outcome evolutions of switchers and stayers with the same period-one treatment. Then, this estimand requires that there be no value of $D_{1}$ such that only switchers have that value, as assumed in Assumption 4. Assumption 5 requires that there are no quasi-stayers: the treatment of all switchers changes by at least $\kappa>0$.
Define the functions:
$\mu_{0}(d):=E[\Delta Y|S=0,D_{1}=d]\quad\text{ and }\alpha_{0}(s,d):=E\left(\left.\frac{S}{\Delta D}\right|D_{1}=d\right)\frac{1-s}{E(1-S|D_{1}=d)}.$
@@
If Assumptions 1, 2, 4, and 5 hold, then for any functions $(s,d)\mapsto\alpha(s,d)$ and $d\mapsto\mu(d)$, respectively defined on $\{0,1\}\times\mathcal{D}_{1}$ and $\mathcal{D}_{1}$, such that $\alpha$ is bounded and $E[|\mu(D_{1})|]<\infty$:
$\delta_{1}=$ $E\left(\frac{1}{E[S]}\left(\frac{S}{\Delta D}-\alpha(S,D_{1})\right)\left[\Delta Y-\mu_{0}(D_{1})\right]\right)$
$=$ $E\left(\frac{1}{E[S]}\left(\frac{S}{\Delta D}-\alpha_{0}(S,D_{1})\right)\left[\Delta Y-\mu(D_{1})\right]\right).$
It readily follows from Lemma 1 and the law of iterated expectations that $\delta_{1}$ is identified by
$E\left[\left(\frac{\Delta Y-E[\Delta Y|S=0,D_{1}]}{\Delta D}\right)\middle|S=1\right]=E\left[\left(\frac{\Delta Y-\mu_{0}(D_{1})}{\Delta D}\right)\middle|S=1\right],$
where the numerator of the ratio inside the expectation is a DID comparing the outcome evolutions of switchers and stayers with the same $D_{1}$. Theorem 1 implies that a related estimand,
$E\left(\frac{1}{E[S]}\left(\frac{S}{\Delta D}-\alpha_{0}(S,D_{1})\right)\left[\Delta Y-\mu_{0}(D_{1})\right]\right),$
also identifies $\delta_{1}$, and is doubly-robust in the functions $\mu_{0}(D_{1})$ and $\alpha_{0}(S,D_{1})$. [^4] We are grateful to Andres Santos for pointing out to us that the AS is identified by a doubly-robust estimand.
If there are quasi-stayers, the AS is still identified, but estimation may be challenging. For any $\eta>0$, let $S_{\eta}=1\{|\Delta D|>\eta\}$ be an indicator for switchers whose treatment changes by at least $\eta$ from period one to two.
@@
$\Delta Y-E(\Delta Y|D_{1},S=0)$ $=Y_{2}(D_{2})-Y_{2}(D_{1})+\Delta Y(D_{1})-E(\Delta Y(D_{1})|D_{1},S=0)$
@@
Therefore, we need to trim quasi-stayers from the estimand in Theorem 1, and let the trimming go to 0, as in Graham and Powell (2012) who consider a related estimand with some quasi-stayers. Accordingly, with quasi-stayers the AS is irregularly identified by a limiting estimand. Then, we conjecture that the AS cannot be estimated at the $\sqrt{n}-$rate, as in Graham and Powell (2012), and as is often the case with target parameters identified by limiting estimands. Therefore, we do not consider estimation of the AS with quasi stayers.
@@ 3.3 Estimation and inference
Let
$g_{0}(d):=E[S/\Delta D|D_{1}=d]\quad\text{and }p_{0}(d):=E[1-S|D_{1}=d].$
It follows from Theorem 1 that estimating $\delta_{1}$ requires estimating the nuisance functions $g_{0}(d)$, $p_{0}(d)$, and $\mu_{0}(d)$ in a first step, before averaging the variable
$\left(\frac{S_{i}}{\Delta D_{i}}-\frac{\hat{g}(D_{i,1})}{\hat{p}(D_{i,1})}(1-S_{i})\right)\left(\Delta Y_{i}-\hat{\mu}(D_{i,1})\right).$
We propose to use cross-fitting to perform this two-step estimation. For any set $A$, let #$\#A$ denote its cardinality. Let the sample be split randomly into two sub-samples, $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$, and let $I_{k}=\#\mathcal{I}_{k}$ for $k=1,2$. We let $\hat{g}^{(2)}$, $\hat{p}^{(2)}$ and $\hat{\mu}^{(2)}$ be nonparametric estimators of $g_{0}$, $p_{0}$ and $\mu_{0}$ based on $\mathcal{I}_{2}$. Then the estimator $\hat{\delta}_{1,\mathsf{DR}}^{(1)}$ is computed in $\mathcal{I}_{1}$ as follows:
$\hat{\delta}_{1,\mathsf{DR}}^{(1)}=\frac{1}{n_{S}^{(1)}}\sum_{i\in\mathcal{I}_{1}}\left(\frac{S_{i}}{\Delta D_{i}}-\frac{\hat{g}^{(2)}(D_{i,1})}{\hat{p}^{(2)}(D_{i,1})}(1-S_{i})\right)\left(\Delta Y_{i}-\hat{\mu}^{(2)}(D_{i,1})\right),$
where $n_{S}^{(1)}:=\sum_{i\in\mathcal{I}_{1}}S_{i}$ and we use the convention that $\hat{\delta}_{1,\mathsf{DR}}^{(1)}=0$ if $n_{S}^{(1)}=0$. The estimator $\hat{\delta}_{1,\mathsf{DR}}^{(2)}$ is computed similarly, permuting the roles of $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$. Finally the cross-fitting doubly-robust estimator is computed as:
$\hat{\delta}_{1,\mathsf{DR}}=\frac{n_{S}^{(1)}}{n_{S}^{(1)}+n_{S}^{(2)}}\hat{\delta}_{1,\mathsf{DR}}^{(1)}+\frac{n_{S}^{(2)}}{n_{S}^{(1)}+n_{S}^{(2)}}\hat{\delta}_{1,\mathsf{DR}}^{(2)}.$
We impose the following condition:
@@
$\mathcal{I}_{1}\perp\!\!\!\perp(D_{i,2},Y_{i,1},Y_{i,2})_{i=1,...,n}|(D_{i,1})_{i=1,...,n}$ and $I_{1}/n\to\pi\in(0,1)$.
We also require that the estimators of the nonparametric nuisance functions satisfy some convergence-rate requirements. For any function $h(\cdot)$, $\hat{h}^{(1)}(.),\hat{h}^{(2)}(.)$ estimators of $h(.)$ based on subsample $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ respectively and $k\in\{1,2\}$, define:
$\left\|\hat{h}-h\right\|_{2,\mathcal{I}_{k}}:=\left(\frac{1}{I_{k}}\sum_{i\in\mathcal{I}_{k}}(\hat{h}^{(3-k)}(D_{1,i})-h(D_{1,i}))^{2}\right)^{1/2},\quad\left\|\hat{h}-h\right\|_{k,\infty}:=\sup_{d\in\mathcal{D}_{1}}\left|\hat{h}^{(3-k)}(d)-h(d)\right|.$
Assumption 7
For $k=1,2$,
  1. (a)
    $\left\|\hat{g}-g_{0}\right\|_{2,\mathcal{I}_{k}}=o_{P}(1)$, $\left\|\hat{\mu}-\mu_{0}\right\|_{2,\mathcal{I}_{k}}=o_{P}(1)$, $\left\|\hat{p}-p_{0}\right\|_{k,\infty}=o_{P}(1)$.
  2. (b)
    $\sqrt{n}\left\|\hat{g}-g_{0}\right\|_{2,\mathcal{I}_{k}}\times\left\|\hat{\mu}-\mu_{0}\right\|_{2,\mathcal{I}_{k}}=o_{P}(1)$, $\sqrt{n}\left\|\hat{p}-p_{0}\right\|_{2,\mathcal{I}_{k}}\times\left\|\hat{\mu}-\mu_{0}\right\|_{2,\mathcal{I}_{k}}=o_{P}(1)$.
As all nuisance functions are conditional expectations with respect to a single variable, these convergence-rate requirements are satisfied by classical nonparametric estimators. For instance, one can use series estimators with a polynomial order chosen by cross validation within subsample $\mathcal{I}_{k}$ to estimate $\hat{\mu}^{(k)}(d)$, $\hat{g}^{(k)}(d)$, and $\hat{p}^{(k)}(d)$. Results in Section 4 of Andrews (1991) imply that those estimators satisfy the rate conditions in Assumption 7, as long as the population functions are twice differentiable. Note also that the mean-square convergence requirements can be established under regularity conditions and appropriate choices of tuning parameters for machine learning estimators such as random forests (Biau, 2012; Wager and Athey, 2018; Syrgkanis and Zampetakis, 2020) and neural networks (Chen and White, 1999; Farrell et al., 2021). [^5] We rely on uniform convergence of the propensity score to ensure that the inverse propensity score is consistent. While uniform convergence may be too strong for machine learning estimators, this condition may be relaxed using automatic debiased machine learning, as proposed in Chernozhukov et al. (2024).
@@
If Assumptions 1-7 hold, $\sqrt{n}(\hat{\delta}_{1,\mathsf{DR}}-\delta_{1})\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{1}))$ where
$\psi_{1}:=\frac{1}{E[S]}\left\{\left(\frac{S}{\Delta D}-E\left(\left.\frac{S}{\Delta D}\right|D_{1}\right)\frac{1-S}{E(1-S|D_{1})}\right)\left[\Delta Y-E(\Delta Y|S=0,D_{1})\right]-\delta_{1}S\right\}.$